Find three numbers such that the first is the sum of the second and third, the second is the square of the third, and the sum of the three numbers is a minimum.
I only was able to get these equations but did not know how to solve it without any numbers.
A=B+C
B^2=C
A+B+C=Min

Answer by ankor@dixie-net.com(22740)

Find three numbers such that the first is the sum of the second and third,
A = B+C
the second is the square of the third,
B = C^2
and the sum of the three numbers is a minimum.
A+B+C
Replace A with B+C
B+C + B + C = 2B + 2C
Replace B with C^2
2C^2 + 2C
Assume we are dealing with integers, let C = 1, then B = 1, Then A = 2
A+B+C = 4, would the be the minimum, wouldn't it?